OpenStudy (anonymous):

Can you factor 2x^4-5x^3+53x^2-125x+75 into 2 quadratic expressions?

OpenStudy (anonymous):

\[2x ^{4}-5x ^{3}+53x^2-125x+75\]

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

Well, reducing fully, \(\normalsize\color{blue}{ (x-1)(2x+3)(x^2+25)\LARGE\color{white}{ \rm │ }}\) .

OpenStudy (solomonzelman):

So yes, you will get \(\normalsize\color{blue}{ (x^2+x-3)(x^2+25)~.\LARGE\color{white}{ \rm │ }}\)

OpenStudy (anonymous):

Ok thanks, and then to find the zeros, I use (x-1)(2x-3)(x+5)(x-5)? @SolomonZelman

OpenStudy (solomonzelman):

No, it is x\(\normalsize\color{blue}{ +}\)25, not -.

OpenStudy (anonymous):

x^2+25=0 --> x=±5?

OpenStudy (solomonzelman):

You will get \(\normalsize\color{blue}{ (x-1)(2x-3)(x^2+25)=0\LARGE\color{white}{ \rm │ }}\) ───────────────── \(\normalsize\color{blue}{ (x-1)=0~~~~⇒~~~~x=1\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ (2x-3)=0~~~~⇒~~~~x=3/2\LARGE\color{white}{ \rm │ }}\) ───────────────── or, \(\normalsize\color{blue}{x^2+25=0\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{x^2=-25\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{x=±5i\LARGE\color{white}{ \rm │ }}\)

OpenStudy (anonymous):

Right, I see, the 25 would be negative when you move it over so you have to use an imaginary number to get the square root.

OpenStudy (solomonzelman):

Yes, exactly :)

OpenStudy (anonymous):

OK, thanks :D

OpenStudy (solomonzelman):

Anytime:)